In a parallelogram ABCD, P divides AB in the ratio 2:5 and Q divides DC in the ratio 3:2. If AC and PQ intersect at R , find the ratios AR:RC and PR:RQ
1
Expert's answer
2019-09-20T11:07:21-0400


1)"AB||CD", "PQ" is transversal line, "\angle RPA" and "\angle CQR" are alternate interior angles, so "\angle RPA=\angle CQR"

2)"AB||CD", "AC" is transversal line, "\angle PAR" and "\angle QCR" are alternate interior angles, so "\angle PAR=\angle QCR"

3)"\angle PRA=\pi-\angle RPA-\angle PAR" in "\Delta APR" , "\angle CRQ=\pi-\angle CQR-\angle CRQ" in "\Delta CRQ"

"\angle PRA=\pi-\angle RPA-\angle PAR=\pi-\angle CQR-\angle CRQ=\angle CRQ"

So "\Delta APR \sim \Delta CQR" and we have "\frac{AP}{CQ}=\frac{AR}{CR}=\frac{PR}{QR}"

Since "\frac{AP}{BP}=\frac{2}{5}", we obtain that "AP=2x" and "BP=5x" for some "x"

Similarly we obtain that "DQ=3y" and "QC=2y" for some "y"

So "7x=AP+BP=AB=CD=CQ+QD=5y" , and "y=1.4x"

We have "\frac{AR}{CR}=\frac{PR}{QR}=\frac{AP}{CQ}=\frac{2x}{2y}=\frac{2x}{2\cdot 1.4x}=\frac{5}{7}"

Answer: "\frac{AR}{CR}=\frac{PR}{QR}=\frac{5}{7}"